OpenStudy (anonymous):

Which of the following represents an equation of a line that passes through (2, 9) and (1, 4)? 5x+y=-19 5x+y=9 -5x+y=1 -5x+y=-1

OpenStudy (puzzler7):

Plug both pairs of values in. If they both work, that is your equation.

OpenStudy (anonymous):

So lets take the first equation and plug in (2,9) 5(2)+(9)=19 So voila the first point makes sense. Lets try (1,4) 5(1)+4=9 not 19 so therefore this equation doesn't pass btwn both points

OpenStudy (anonymous):

The equation passes through the first point but not the second. Now looking at the second equation I can see that it wont pass through the first point but only the second

OpenStudy (anonymous):

Now lets look at the last equation -5x+y=-1 Lets plug in (2,9) -5(2)+9=-1 Lets plug in the second point (1,4) -5(1)+4=-1 There we see that the last equation passes those 2 points on the graph

OpenStudy (anonymous):

Thank you so much Sam! I really appreciate the help!

OpenStudy (anonymous):

No problem :D